OpenStudy (selena2345):

Solve The Equation.Then Check The Solution m+3(m-4)=16

OpenStudy (anonymous):

distribute the 3 first hint

OpenStudy (selena2345):

can you show the whole work for the problem ?

OpenStudy (anonymous):

m+3(m-4)=16 m+3m-12=16 4m-12=16 4m=16+12 m=(16+12)/4 m=4+3=7

OpenStudy (ashleyisakitty):

m + 3(-4 + m) = 16 m + (-4 * 3 + m * 3) = 16 m + (-12 + 3m) = 16 Reorder then combine like terms. -12 + 4m = 16 solve for m disturbute -12 + 12 + 4m = 16 + 12 Combine like terms 4m = 16 + 12 4m=28 divide each side by four M=?

OpenStudy (ashleyisakitty):

hope my steps can help you in the future